### v5

# Symmetry & Critical Points for Symmetric Tensor Decomposition Problems

Abstract.
We consider the nonconvex optimization problem associated with the decomposition of a real symmetric tensor into a sum of rank-one terms. Use is made of the rich symmetry structure to construct infinite families of critical points represented by Puiseux series in the problem dimension, and so obtain precise analytic estimates on the objective function value and the Hessian spectrum. The results enable an analytic characterization of various obstructions to local optimization methods, revealing, in particular, a complex array of saddles and minima that differ in their symmetry, structure, and analytic properties. A notable phenomenon, observed for all critical points considered, concerns the index of the Hessian increasing with the objective function value.
1. Introduction
The need to decompose a real symmetric tensor into a sum of rank-one terms arises naturally in various scientific domains [1, 2, 3, 4, 5]. Standard approaches [6, 7, 8, 9, 10, 11] involve tackling a nonconvex optimization problem associated with an order-$n$ symmetric tensor $T$ on $\mathbb{R}^{d}$, given by
$\min_{\stackrel{{\scriptstyle W\in M(k,d),}}{{\boldsymbol{\alpha}\in\mathbb{R}^{k}}}}{\mathcal{L}}(W,\boldsymbol{\alpha})\coloneqq\left\|\sum_{i=1}^{k}\alpha_{i}\mathbf{w}_{i}^{\otimes{}n}-T\right\|^{2}$ (1.1)
for fixed $k\in\mathbb{N}$, where $M(k,d)$ denotes the set of $k\times d$-real matrices, $\mathbf{w}_{i}$ the $i$th row of a weight matrix $W\in M(k,d)$, and $\|\cdot\|$ a tensor norm. It has been observed [10, 11, 12, 13] that these approaches often fail to find valid decompositions in various parameter regimes encountered in practice. This article focuses on the geometric obstructions that arise in the associated optimization landscape, particularly saddles and minima, which contribute to such failures. Rather than striving for general results, we focus on symmetric tensors that are invariant to certain actions of subgroups of $S_{m}$, the symmetric group over $m$ symbols. This choice yields analytically tractable problems whose geometry exposes key obstructions to symmetric tensor decomposition—the aim of this work.
2. Main results
After a quick review of the setting studied and key related concepts, the introduction continues with a detailed description of the main results. We conclude with a brief survey of related work and outline of the structure of the paper. Some familiarity with real algebraic geometry and group theory is assumed, but otherwise, key ideas and concepts are introduced as needed, with the exception of several irreducible representations of $S_{d}$ that are explicitly referred to below (see Section 5) and may be safely skipped without loss of continuity. Proofs are deferred to the appendix.
We consider a natural action of $S_{k}\times S_{d}$ on the parameter space $M(k,d)$ (orthogonal): the first factor permutes rows, the second columns. Given a weight matrix $W\in M(k,d)$, the largest subgroup of $S_{k}\times S_{d}$ fixing $W$ is called the isotropy group and is used as a means of measuring the symmetry of $W$ (see Section 4). For example, for $k=d$, the isotropy group of the identity matrix is the diagonal subgroup $\Delta S_{d}\coloneqq\{(\pi,\pi)~|~\pi\in S_{d}\}\subseteq S_{d}\times S_{d}$. In [14], it was shown that, if $T$ possesses certain invariance properties, $\mathcal{L}$} is $S_{k}\times S_{d}$-invariant. An intricate mechanism, connecting the symmetry of adjacent critical points of invariant functions [15], then dictates the emergence of critical points of $\mathcal{L}$} that exhibit proper isotropy subgroups breaking the symmetry of $S_{k}\times S_{d}$, as is verified through numerical and algebraic methods. In this work, this phenomenon of symmetry breaking is employed to derive sharp analytic results for the optimization landscape associated with $\mathcal{L}$}. More specifically,
  • The invariance properties of a broad class of optimization problems are formalized, encompassing those associated with symmetric tensor decomposition and two-layer ReLU networks. These properties enable a detailed study of symmetric critical points, particularly ones exhibiting large isotropy groups.
  • We construct infinite families of critical points represented by Puiseux series in $d$ and derive sharp analytic estimates of the Hessian spectrum. The tensor norms considered are the Frobenius norm and the cubic-Gauss norm defined in (2.2).
  • The complex array of saddles and minima so revealed exhibits various phenomena affecting the behavior of local optimization methods, as discussed in length in Section 2.2. For example, saddles with a higher objective value are seen to exhibit a larger number of descent directions.
A general introduction to the symmetry-breaking framework is provided in Section 5. For reasons of exposition, the applications in this article are presented primarily for third-order symmetric tensors under the assumption $k=d$, following for example [13]. Moreover, we shall focus on symmetric tensors admitting an orthonormal decomposition, that is, $T=\sum_{i=1}^{d}\mathbf{v}_{i}^{\otimes{}3}$ where $\mathbf{v}_{1},\dots,\mathbf{v}_{d}$ are orthonormal. The methods, however, are quite general. Orthonormal decomposability of $T$ (see related [16]) is not required; other choices of order, dimensionality, rank and target tensors with possibly lesser symmetry may be considered.
2.1. Global minimizers
We begin by providing a simple characterization of the global minimizers of $\mathcal{L}$}. Our study of families of critical points is essentially invariant to orthogonal transformations of the domain of $\mathcal{L}$}, and so throughout we assume (without loss of generality, Section A) that the target tensor is $T_{\mathbf{e}}\coloneqq\sum_{i=1}^{d}\mathbf{e}_{i}^{\otimes{}3}$, where $\mathbf{e}_{i}$ denotes the $i$th unit vector. Moreover, since $n$ is odd, all $\alpha_{i}$ may be absorbed into $\mathbf{w}_{i}$ and are thus set to 1. Given a weight matrix $W\in M(d,d)$, the action of $S_{d}\times S_{d}$, perturbing rows and columns of $W$, typically yields additional points constituting $(S_{d}\times S_{d})W$, the group orbit of $W$. For our choice $T=T_{\mathbf{e}}$, the objective loss $\mathcal{L}$} is $S_{d}\times S_{d}$-invariant (see Section 4) and so points on the orbit of a critical point are also critical and share the same loss and Hessian spectrum, see generally [17]. Since ${\mathcal{L}}\geq 0$ and ${\mathcal{L}}(I_{d})=0$, the identity matrix $I_{d}$, as well as all points lying on $(S_{d}\times S_{d})I_{d}$, are global minimizers. That the converse holds is a consequence of Kruskal’s criterion for identifiability [18].
Proposition 1.
Independently of the choice of the tensor norm, a point $W\in M(d,d)$ is a global minimizer of $\mathcal{L}$} iff $W\in(S_{d}\times S_{d})I_{d}$, the set of all $d\times d$ permutation matrices. In particular, the set of global minimizers and the set of points defining valid decompositions for $T_{\mathbf{e}}$ are identical.
Versions of the proposition involving the action of the Hyperoctahedral group hold for tensors of order $n\geq 4$ and may be proven by an extension to Kruskal’s criterion given in [19] (see [20, Prop. 4.14] for an analog result for ReLU).
By Proposition 1, issues concerning the existence of global solutions for (1.1) do not arise [21]; a global solution always exists. Moreover, the number of global minimizers is equal to the length of the orbit $|(S_{d}\times S_{d})I_{d}|=d!$. The large number of global minimizers accounts for, in part, the complex array of critical points occurring for $\mathcal{L}$}. In the general case, the orbit-stabilizer theorem states that the smaller the isotropy group, the larger the group orbit. Below we indicate orbit lengths at appropriate points.
Although the set of global minimizers is discrete, this need not be the case for other families of critical points. For example, a continuous of family of critical points, studied in some depth below, is given by $\mathfrak{C}_{5,t}(d)\coloneqq I_{d-2}\oplus\begin{bmatrix}(1-t^{3})^{1/3}&0\\ t&0\end{bmatrix}$.
Proposition 2.
For all $t\in\mathbb{R}$ and $d\geq 2$, $\mathfrak{C}_{5,t}(d)$ is a critical point having ${\mathcal{L}}(\mathfrak{C}_{5,t}(d))=1$. In particular, the level set ${\mathcal{L}}^{-1}(\{1\})$ is not bounded and so $\mathcal{L}$} is not a proper map.
The family of critical points $\mathfrak{C}_{5,0}$ in fact defines saddles. The proof is based on the curve selection (CSL) lemma, used in the framework of tangency arcs; see [15].
2.2. Symmetric Saddles
Contrarily to global minimizers, the set of saddle points exhibits a highly complex and intricate structure. Below, we consider two tensor norms widely used in the literature (see Section 4): the Frobenius norm $\|\cdot\|_{F}$ and the cubic-Gaussian norm $\|\cdot\|_{\mathcal{N}}$ defined by
$\|S-T\|_{\mathcal{N}}\coloneqq\mathbb{E}_{\mathbf{x}\sim\mathcal{N}(0,I_{d})}[\langle S-T,\mathbf{x}^{\otimes{}3}\rangle^{2}],$ (2.2)
with $\mathcal{N}(0,I_{d})$ denoting the standard multivariate Gaussian distribution, see [14, 22] and references therein. The objective function will be respectively denoted by ${\mathcal{L}}_{F}$ and ${\mathcal{L}}_{\mathcal{N}}$. A consequence of the study of the two different norms is that phenomena which may otherwise be deemed norm-specific can be examined and verified in separate settings, see for example (2.3) below.
The Frobenius norm
We begin by considering families of critical points $\{W(d)\in M(d,d)\}_{d\in\mathbb{N}}$ occurring for the Frobenius norm. The entries of each family are expressed in terms of Puiseux series in $d$, giving a critical point of ${\mathcal{L}}_{F}:M(d,d)\to\mathbb{R}$ for every integer $d\in\mathbb{N}$.
Theorem 1.
The sequences of weight matrices $\mathfrak{C}_{i}(d)\in M(d,d),~i\in\{I,0,1,2,3,4,5\}$, given in Section 5, define families of critical points of ${\mathcal{L}}_{F}$ for the target tensor $T_{\mathbf{e}}$. The loss, the Hessian spectrum and the orbit length are given in Table 1 and Table 2.
$\mathfrak{C}_{0}$ $\mathfrak{C}_{1}$ $\mathfrak{C}_{2}$ $\mathfrak{C}_{I}$
Loss $d$ d1/dd-1/d d1/dd-1/d 0
Isotropy $S_{d}\times S_{d}$ $S_{d}\times S_{d}$ $\Delta S_{d}$ $\Delta S_{d}$
Orbit length 11 11 $d!$ $d!$
Comp. Multiplicity
$\mathfrak{t}$ 11 0 $18/d$ $18/d$ 1818
$12/d$ 66
$\mathfrak{s}_{d}$ d1d-1 0 0 $12/d$ 1818
0 6/d-6/d 0 66
6/d-6/d 66
$\mathfrak{s}_{d}\otimes{}\mathfrak{s}_{d}$ $(d-1)^{2}$ 0 12/d-12/d
$\mathfrak{x}_{d}$ (d1)(d2)/2(d-1)(d-2)/2 12/d-12/d 66
$\mathfrak{y}_{d}$ d(d3)/2d(d-3)/2 12/d-12/d 66
Table 1. The loss and the Hessian spectrum given in $o(d^{-1})$-terms for the families of critical points $\mathfrak{C}_{1}$ and $\mathfrak{C}_{2}$, and in precise terms for $\mathfrak{C}_{I}$ giving the identity matrices (global minimizers, see Proposition 1) and for $\mathfrak{C}_{0}$ the zero matrices. The families $\mathfrak{C}_{1}$ and $\mathfrak{C}_{2}$ define saddles for sufficiently large $d$. The family $\mathfrak{C}_{0}$ (as well as $\mathfrak{C}_{5}$ in Table 2 below) also defines saddles, but due to the singularity of the Hessian, the proof requires going beyond second-order analysis; see [15]. The orbit length indicates the number of different families of critical points that occur by the action of $S_{d}\times S_{d}$, having in particular identical loss and Hessian spectrum which is listed by the respective irreducible representation of $S_{d}$, Section 5.
$\mathfrak{C}_{3}$ $\mathfrak{C}_{4}$ $\mathfrak{C}_{5}\coloneqq\mathfrak{C}_{5,0}$ Isotropy $\Delta(S_{d-1}\times S_{1})$ $\Delta(S_{d-2}\times S_{2})$ $\Delta(S_{d-2}\times S_{1}^{2})$ Loss $\quad\quad d-1$ $3/2$ 11 Orbit length $\quad d\cdot d!$ d(d1)d!/2d(d-1)d!/2 d(d1)d!d(d-1)d! Eigenvalue #\# Eigenvalue #\# Eigenvalue #\# 1818 11 1818 d2d-2 1818 d1d-1 66 d1d-1 99 11 66 $(d-1)^{2}$ $18/d$ 1 66 $d^{2}-3d+2$ 0 $d$ $12/d$ d1d-1 33 d2d-2 $6/d$ 1 0 d1d-1 0 2d42d-4 3-3 11 6/d-6/d d2d-2 6-6 11 12/d-12/d $d^{2}-5d+5$
Table 2. The loss and the Hessian spectrum given in $o(d^{-1})$-terms for the family of critical points $\mathfrak{C}_{3}$ and in precise terms for the families of critical points $\mathfrak{C}_{4}$ and $\mathfrak{C}_{5}$. Similarly to $\mathfrak{C}_{0}$ appearing in Table 1, it is not possible to determine whether $\mathfrak{C}_{5}$ defines a family of saddles or minima by inspecting the Hessian spectrum alone as the Hessian is positive semi-definite and singular. In [15], the existence of real analytic descent curves for $\mathfrak{C}_{5}$ is proved using the CSL [23], showing in particular that $\mathfrak{C}_{5}$ defines saddles.
In [24], several families of minima of ${\mathcal{L}}_{F}$ are described. However, only the $k=1$ case is considered, and the Hessian spectrum is provided only for certain local minima and with limited accuracy. In contrast, the symmetry breaking approach applies in principle to any choice of $k$ and $d$, and yields an analytic description of the Hessian spectrum (as well as the critical value) at both local minima and saddle points. We note that although the Hessian has $d^{2}$ eigenvalues, the number of distinct eigenvalues is fixed and independent of the dimension; a consequence of the symmetry structure, namely, the multiplicity of irreducible representations occurring in the respective isotypic decomposition being independent of the dimension (Section 5).
The saddle points given in Theorem 1 behave rather differently in terms of the dependence of their objective value and Hessian spectrum on the dimension, cf. [25]. While the Hessian spectrum at $\mathfrak{C}_{0},\mathfrak{C}_{1}$ and $\mathfrak{C}_{2}$ converges to zero, some of the positive eigenvalues of $\mathfrak{C}_{3}$ and $\mathfrak{C}_{4}$ are bounded away from zero. In fact, for $\mathfrak{C}_{4}$, except for two negative eigenvalues, $\Theta(d^{2})$ of the eigenvalues are strictly positive, and so despite being saddles, critical points belonging to $\mathfrak{C}_{4}$ may be hard to escape via gradient-based methods. In contrast, at $\mathfrak{C}_{3}$, $\Theta(d^{2})$ of the Hessian eigenvalues are negative (of the order of $1/d$). Comparing the objective value, $\mathcal{L}$$_{\mathcal{F}}$ evaluated at either $\mathfrak{C}_{0},\mathfrak{C}_{1},\mathfrak{C}_{2}$ or $\mathfrak{C}_{3}$ scales as $\Theta(d)$. However, critical values given by saddles need not grow linearly with $d$. For example, at $\mathfrak{C}_{4}$, the critical value is fixed and equal to $3/2$.
Our findings suggest a direct connection between the function value at non-zero critical points and the index, that is, the number of negative Hessian eigenvalues, at a (not necessarily hyperbolic) critical point: the larger the critical value, the larger the index—a desirable property for successful use of local optimization methods. Quantitatively, for a general family critical points $\mathfrak{C}(d)$, this may be expressed in the form
$\frac{{\mathcal{L}}(\mathfrak{C}(d))}{d}\approx\frac{\mathrm{Index}(\mathfrak{C}(d))}{d^{2}},$ (2.3)
or, rather, in a more general form replacing second-order descent directions, counted by $\mathrm{Index}(\cdot)$, with ones corresponding to second- or higher-order derivatives. The above phenomenon (indeed, its quantitative version (2.3)) also occurs for the cubic-Gaussian norm (Table 3 below) and for two-layer ReLU networks [25] (cf. [26, Remark 2] on the expected initial value growing linearly with $d$ under Xavier initialization). It is therefore natural to conjecture that the above, or a qualitatively similar, relation between the function value at critical points and the index applies more broadly to non-convex functions exhibiting symmetry breaking phenomena of the nature considered in this work.
The cubic-Gaussian norm
Although the geometry of the associated optimization landscape depends on the choice of the tensor norm, by Proposition 1, the set of global minimizers giving valid tensor decomposition is identical for all norms. The critical set, however, differs quite significantly. We illustrate this for the cubic-Gaussian norm (2.2).
Theorem 2.
The sequences of weight matrices $\mathfrak{D}_{I}(d),\mathfrak{D}_{1}$ and $\mathfrak{D}_{2}$, described in Section 5, define families of critical points of ${\mathcal{L}}_{\mathcal{N}}$ for the target tensor $T_{\mathbf{e}}$. The loss, the Hessian spectrum and the orbit length are given in Table 3.
The loss at $\mathfrak{D}_{I},\mathfrak{D}_{1}$ and $\mathfrak{D}_{2}$ scale similarly to ones seen for the Frobenius norm. However, the Hessian spectrum behaves rather differently. For example, the Hessian spectrum at the families of minima $\mathfrak{D}_{I}$ is extremely skewed having $\Theta(d^{2})$ positive eigenvalues concentrated near zero and $\Theta(d)$ eigenvalues growing linearly with $d$. In addition, modulo $\Theta(d)$ eigenvalues, the spectrum of the families of third-order saddles $\mathfrak{D}_{2}$ and $\mathfrak{D}_{I}$ essentially agree. The two phenomena: (a) highly-skewed Hessian spectra (b) largely agreeing at certain critical points, have also been observed for two-layer ReLU networks [27, 25]. Lastly, although $\mathfrak{C}_{I}$ and $\mathfrak{D}_{I}$ represent the same family of global minima, the Hessian spectra differ significantly: the former is fixed, the latter depends on $d$. This further emphasizes the dependence of the geometry of the associated optimization landscape on the choice of the tensor norm.
$\mathfrak{D}_{1}$ $\mathfrak{D}_{I}$ $\mathfrak{D}_{2}$
Isotropy $S_{d}\times S_{d}$ $\Delta S_{d}$ $\Delta(S_{d-1}\times S_{1})$
Loss $\frac{48d}{5}-\frac{36}{5}-\frac{12}{5d}$ 0 $6+18/d$
Multiplicity 11 $d!$ $d\cdot d!$
Comp. Degree
$\mathfrak{t}$ 1 $\frac{162\cdot{75}^{\frac{1}{3}}}{5}{d}^{\frac{1}{3}}+\frac{144\cdot{75}^{\frac{1}{3}}}{5}d^{\frac{-2}{3}}$ 108 0
$18d+180$ 0
108
$18d+18$
$18d+162$
$\mathfrak{s}_{d}$ d1d-1 0 108
$\frac{18\cdot{75}^{\frac{1}{3}}}{5}{d}^{\frac{1}{3}}-\frac{32\cdot{75}^{\frac{1}{3}}}{5}d^{\frac{-2}{3}}$ 36
$18d+180$
$\mathfrak{s}_{d-1}$ d2d-2 0
36
36
108
$18d+162$
$\mathfrak{s}_{d}\otimes{}\mathfrak{s}_{d}$ $(d-1)^{2}$ $\frac{72\cdot{75}^{\frac{1}{3}}}{25}{d}^{\frac{1}{3}}-\frac{304\cdot{75}^{\frac{1}{3}}}{25}d^{\frac{-2}{3}}$
$\mathfrak{x}_{d}$ (d1)(d2)/2(d-1)(d-2)/2 36
$\mathfrak{x}_{d-1}$ (d2)(d3)/2(d-2)(d-3)/2 36
$\mathfrak{y}_{d}$ d(d3)/2d(d-3)/2 36
$\mathfrak{y}_{d-1}$ (d1)(d4)/2(d-1)(d-4)/2 36
Table 3. The loss and the Hessian spectrum of three families of critical points given in Section 5.1, with $\mathfrak{D}_{I}$ defining global solutions and $\mathfrak{D}_{1}$ saddles. The Hessian spectrum is given modulo $o(d^{-\frac{2}{3}})$-terms and is listed by the respective irreducible representation of $S_{d}$ (Section 5).
2.3. Related work
We relate our results to the existing literature.
Symmetric tensor decomposition.
The tensor decomposition problem considered in this work, also referred to as the polyadic decomposition, was introduced under the names CANDECOMP and PARAFAC by respectively [6] and [28]. Standard approaches involving nonconvex optimization formulation include [6, 7, 8, 9, 10, 11]. See [12] for a survey of methods and applications. As noted in the introductory comments, it has been observed [10, 11, 12, 13] that these approaches have a relatively high likelihood of failure across various parameter regimes encountered in practice. Theoretical guarantees are available in some cases [24, 29, 30], but none of the latter directly addresses the optimization landscape for $k>1$. A computationally intensive yet provable method, assuming $T$ has rank $O(n^{\frac{d-1}{2}})$, was presented in [31]. For $n=3$ and $k\leq d$, a theoretically convenient method was derived based on simultaneous diagonalization of matrix slices [7]. For $k=4$ and $k=O(n^{2})$, the work [32] introduced a provable algorithm using matrix eigendecompositions, which was later robustified using ideas from the sums-of-squares hierarchy in [33]. In [34], a tensor power method, constructed from a matrix flattening of $T$, was applied to sequentially extract the components $\mathbf{w}_{i}$.
2.4. Organization of the paper.
In Section 3, we briefly review relevant background from multilinear algebra. In Section 4, the invariance properties of a large class of objective functions, subsuming $\mathcal{L}$}, are studied. In Section 5, the framework of symmetry breaking is introduced, and Puiseux series are constructed for various families of critical points using Newton polyhedra.
3. Preliminaries
A formal discussion of our results requires some familiarity with multilinear algebra. Parts of the exposition follow [14].
A real tensor of order $n$ is an element of the tensor product of $n$ vector spaces $E_{1}\otimes\ldots\otimes E_{n}$, where $E_{i}$ are real vector spaces. Upon choosing bases for each factor $E_{i}$, we may identify a tensor with a multi-dimensional array in $\mathbb{R}^{d_{1}\times\ldots\times d_{n}}$ with $d_{i}=\dim(E_{i})$. We write $T_{i_{1},\ldots,i_{n}}$ for the $(i_{1},\ldots,i_{n})$’th coordinate of $T$. Given vectors $\mathbf{v}_{i}\in\mathbb{R}^{d_{i}}$, $i=1,\ldots,n$, we write $\mathbf{v}_{1}\otimes\ldots\otimes\mathbf{v}_{n}$ for the outer product, namely the element in $\mathbb{R}^{d_{1}}\otimes\ldots\otimes\mathbb{R}^{d_{n}}$ such that $(\mathbf{v}_{1}\otimes\ldots\otimes\mathbf{v}_{n})_{i_{1},\ldots,i_{n}}=({v_{1}})_{i_{1}}\ldots(v_{n})_{i_{n}}$, and $\mathbf{v}^{\otimes n}:=\mathbf{v}\otimes\ldots\otimes\mathbf{v}$ for $n$-times products. The Frobenius inner product of two tensors $T$ and $S$ of the same shape is defined by $\langle T,S\rangle_{F}=\sum_{i_{1},\ldots,i_{n}}T_{i_{1}\ldots i_{n}}S_{i_{1}\ldots i_{n}}$. It is easy to verify that for the Frobenius inner product
$\langle\mathbf{v}_{1}\otimes\ldots\otimes\mathbf{v}_{n},\mathbf{w}_{1}\otimes\ldots\otimes\mathbf{w}_{n}\rangle_{F}=\prod_{i=1}^{n}\langle\mathbf{v}_{i},\mathbf{w}_{i}\rangle$ (3.4)
for all $\mathbf{v}_{i},\mathbf{w}_{i}\in\mathbb{R}^{d_{i}}$. In particular, $\langle\mathbf{v}^{\otimes n},\mathbf{w}^{\otimes n}\rangle_{F}=\langle\mathbf{v},\mathbf{w}\rangle^{n}$.
A tensor $T\in(\mathbb{R}^{d})^{\otimes{}n}$ is symmetric if it is invariant under permutation of indices, that is, if $T_{i_{1},\ldots,i_{n}}=T_{i_{\sigma(1)},\ldots,i_{\sigma(n)}}$ for any permutation $\sigma\in S_{n}$. The space of symmetric tensors of order $n$ on $\mathbb{R}^{d}$, denoted by ${\rm\mathrm{Sym}}^{n}(\mathbb{R}^{d})$, is $\binom{d+n-1}{n}-$dimensional and is isomorphic to the space of homogeneous polynomials of degree $n$ in $d$ variables. A natural isomorphism is given by the map
$S\mapsto P_{S},~~P_{S}(x)\coloneqq\langle S,x^{\otimes r}\rangle_{F}.$ (3.5)
When $n=2$, this is the usual correspondence between symmetric matrices and quadratic forms (see [1, Section 3.1] for more details). The two norms considered in this work are induced from an inner product over ${\rm\mathrm{Sym}}^{n}(\mathbb{R}^{d})$: the Frobenius inner product $\langle\cdot,\cdot\rangle_{F}$, and
$\langle S,T\rangle_{\mathcal{D}}\coloneqq\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\langle S,\mathbf{x}^{\otimes n}\rangle_{F}\langle T,\mathbf{x}^{\otimes n}\rangle_{F}],$ (3.6)
where $\mathcal{D}$ is a distribution on $\mathbb{R}^{d}$ equivalent to Lebesgue measure, chosen so that the above expectation is finite for any $S,T\in\mathrm{Sym}^{n}(\mathbb{R}^{d})$.
Fact 1 ([14], Notation as above).
The bivariate function $\langle\cdot,\cdot\rangle_{\mathcal{D}}:\mathrm{Sym}^{n}(\mathbb{R}^{d})\times\mathrm{Sym}^{n}(\mathbb{R}^{d})\to\mathbb{R}$ defines an inner product on $\mathrm{Sym}^{n}(\mathbb{R}^{d})$.
Note that $\langle\cdot,\cdot\rangle_{\mathcal{D}}$ is not positive definite over $\mathbb{R}^{\otimes{}d}$ and so the explicit restriction to $\mathrm{Sym}^{n}(\mathbb{R}^{d})$ is necessary. Moreover, in terms of the coefficients of $P_{S}(x)=\sum a_{p_{1},\ldots,p_{d}}x_{1}^{p_{1}}\ldots x_{d}^{p_{d}}$ and $P_{T}(x)=\sum b_{p_{1},\ldots,p_{d}}x_{1}^{p_{1}}\ldots x_{d}^{p_{d}}$, the Frobenius inner product on $\mathrm{Sym}^{n}(\mathbb{R}^{d})$ is given by
$\langle S,T\rangle_{F}=\sum_{p_{1}+\ldots+p_{d}=n}\binom{n}{p_{1},\ldots,p_{d}}a_{p_{1},\ldots,p_{d}}b_{p_{1},\ldots,p_{d}},$ (3.7)
and the inner product $\langle\cdot,\cdot\rangle_{\mathcal{D}}$ reads
$\langle T,S\rangle_{\mathcal{D}}=\sum_{\begin{subarray}{c}p_{1}+\ldots+p_{d}=n\\ q_{1}+\ldots+q_{d}=n\end{subarray}}\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[x_{1}^{p_{1}+q_{1}}\ldots x_{d}^{p_{d}+q_{d}}]a_{p_{1},\ldots,p_{d}}b_{q_{1},\ldots,q_{d}}.$ (3.8)
For the standard multivariate normal distribution $\mathcal{D}=\mathcal{N}(0,I_{d})$, the above may be given explicitly,
$\langle S,T\rangle_{\mathcal{N}}=\sum h_{\mathbf{p}+\mathbf{q}}a_{p_{1},\ldots,p_{d}}b_{p_{1},\ldots,p_{d}},$ (3.9)
$\mbox{ where }h_{\mathbf{r}}:=\begin{cases}0&\mbox{if }r_{i}\mbox{ is odd for some }i\in\{1,\ldots,d\},\\ \prod_{i=1}^{d}{(r_{i}-1)!!}&\mbox{otherwise. }\end{cases}$
Remark 1.
As noted in [14], comparing (3.7) and (3.8), it is seen that no data distribution induces the Frobenius inner product. Indeed, in (3.8), whether a term $a_{p_{1},\ldots,p_{d}}$ is multiplied by $b_{q_{1},\ldots,q_{d}}$ depends only on ${p_{1}+q_{1}},\allowbreak\ldots,{p_{d}+q_{d}}$, and so in particular if the coefficient of $a_{p_{1},\ldots,p_{d}}b_{q_{1},\ldots,q_{d}}$ is non-zero then, assuming, e.g., $p_{1}>0$, so is the coefficient corresponding to $a_{p_{1}-1,\ldots,p_{d}}b_{q_{1}+1,\ldots,q_{d}}$. In contrast, the coefficient of $a_{p_{1},\ldots,p_{d}}b_{q_{1},\ldots,q_{d}}$ in the Frobenius product is non-zero if and only if $(p_{1},\ldots,p_{d})=(q_{1},\ldots,q_{d})$.
A symmetric tensor $T\in\mathrm{Sym}^{n}(\mathbb{R}^{d})$ is said to be of (real) rank one if $T=\alpha\mathbf{w}^{\otimes n}$ for some $\alpha\in\mathbb{R}\setminus\{0\}$ and $\mathbf{w}\in\mathbb{R}^{d}\setminus\{0\}$. A tensor is (real symmetric) rank $k$ if it can be written as a linear combination of $k$ rank-one tensors $T=\alpha_{1}\mathbf{w}_{1}^{\otimes n}+\ldots+\alpha_{k}\mathbf{w}_{k}^{\otimes n}$, but not as a combination of k1k-1 rank-one tensors. For $n=2$, this definition agrees with the usual notion of rank of symmetric matrices.
4. Invariance Properties
Assume the target tensor $T$ defining (1.1) is given by $T=\sum_{i=1}^{h}\beta_{i}\mathbf{v}_{i}^{\otimes n}$ for some $V\in M(h,d)$, $\boldsymbol{\beta}\in\mathbb{R}^{h}$ and $h\in\mathbb{N}$. Making the dependence of $\mathcal{L}$} on $T$ explicit, it follows by Proposition 1 that the task of finding a rank $k$ decomposition may be equivalently stated as that of finding a global minimizer for
${\mathcal{L}}(W,\boldsymbol{\alpha};V,\boldsymbol{\beta})=\left\|\sum_{i=1}^{k}\alpha_{i}\mathbf{w}_{i}^{\otimes n}-\sum_{i=1}^{h}\beta_{i}\mathbf{v}_{i}^{\otimes n}\right\|^{2}$ (4.10)
under some tensor norm (in this work, the Frobenius and the cubic-Gaussian norm). Pairs of $(W,\boldsymbol{\alpha})$ with ${\mathcal{L}}(W,\boldsymbol{\alpha};V,\boldsymbol{\beta})=0$ therefore give different rank $k$ decompositions. The extended formulation of the loss function allows a natural description of invariance properties of $\mathcal{L}$} that follow by non-inherent symmetries of $T$ (namely, ones occurring beyond those that are given by the mere definition of symmetric tensors).
Remark 2.
If the rank of $T$ is exactly $k$, the above problem reduces to that of Waring decompositions by the identification given in (3.5) [35].
It will be convenient to adopt a general measure of similarity between vectors in $\mathbb{R}^{d}$, here referred to as a kernel. The use of kernels allows a succinct description of the assumptions used along the derivation of the invariance properties of $\mathcal{L}$}, and yields general results, subsuming in particular settings considered in this work, with no additional cost. The choice to express invariance properties in terms of kernels may be motivated as follows. Observe that
${\mathcal{L}}(W,\boldsymbol{\alpha};V,\boldsymbol{\beta})=\langle\sum_{i=1}^{d}\alpha_{i}\mathbf{w}_{i}^{\otimes n}-\sum_{i=1}^{h}\beta_{i}\mathbf{v}_{i}^{\otimes n},\sum_{i=1}^{d}\alpha_{i}\mathbf{w}_{i}^{\otimes n}-\sum_{i=1}^{h}\beta_{i}\mathbf{v}_{i}^{\otimes n}\rangle_{F}$
$=\sum_{i,j=1}^{k}\alpha_{i}\alpha_{j}\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{n}-2\sum_{i=1}^{k}\sum_{j=1}^{h}\alpha_{i}\beta_{j}\langle\mathbf{w}_{i},\mathbf{v}_{j}\rangle^{n}+\sum_{i,j=1}^{h}\beta_{i}\beta_{j}\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle^{n}.$
We may replace $\langle\cdot,\cdot\rangle^{n}$ by a general kernel function ${\kappa}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ satisfying ${\kappa}(\mathbf{w},\mathbf{v})={\kappa}(\mathbf{v},\mathbf{w})$ and so define,
${\mathcal{L}}_{\kappa}$ $(W,\boldsymbol{\alpha};V,\boldsymbol{\beta})\coloneqq$ (4.11)
$\sum_{i,j=1}^{k}\alpha_{i}\alpha_{j}{\kappa}(\mathbf{w}_{i},\mathbf{w}_{j})-2\sum_{i=1}^{k}\sum_{j=1}^{h}\alpha_{i}\beta_{j}{\kappa}(\mathbf{w}_{i},\mathbf{v}_{j})+\sum_{i,j=1}^{h}\alpha_{i}\beta_{j}{\kappa}(\mathbf{v}_{i},\mathbf{v}_{j}).$
We note that in contrast to the conventional use of kernels in machine learning, in (4.11), $\mathbf{w}_{i}$ are optimization variables. That is, both $\boldsymbol{\alpha}$ and $W$ may be adjusted during the optimization process. Moreover, for the purpose of establishing the invariance properties of $\mathcal{L}$}, $\kappa$} is not required to be positive definite.
We shall be interested in a class of distribution-dependent kernels given by
${\kappa}_{\mathcal{D}}(\mathbf{w},\mathbf{v})=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\rho(\langle\mathbf{w},\mathbf{x}\rangle)\rho(\langle\mathbf{v},\mathbf{x}\rangle)],$ (4.12)
with $\mathcal{D}$ denoting a distribution over $\mathbb{R}^{d}$ and $\rho:\mathbb{R}\to\mathbb{R}$ a measurable function. For example, the choice of $\rho(x)=\max\{x,0\}$, the ReLU activation function, and $\mathcal{D}=\mathcal{N}(0,I_{d})$, is used in the study of two-layer ReLU networks [36].
Remark 3.
If the kernel function $\kappa$} is positive definite, as ones considered in this work, ${\mathcal{L}}_{\kappa}$ can be expressed in terms of inner products of the rows of $W$ and $V$ suitably mapped into the (unique) reproducing kernel Hilbert space corresponding to $\kappa$} [37].
Fact 2 ([14], Notation as above).
For the tensor norm (3.6), ${\mathcal{L}}\equiv{\mathcal{L}}_{{\kappa}_{\mathcal{D}}}$, the latter given by (4.11) with kernel function (4.12) using $\rho(x)=x^{n}$.
The data-dependent kernel corresponding to $\mathcal{D}=\mathcal{N}(0,I_{d})$ and $\rho(x)=x^{3}$ is referred to by the cubic-Gaussian kernel [14], and is given by
${\kappa}_{\mathcal{N}}(\mathbf{w},\mathbf{v})$ $={6}\langle\mathbf{w},\mathbf{v}\rangle_{F}^{3}+9\|\mathbf{w}\|_{2}^{2}\|\mathbf{v}\|_{2}^{2}\langle\mathbf{v},\mathbf{w}\rangle_{F}.$ (4.13)
Definition 1.
We say that a kernel $\kappa$} is permutation-invariant if ${\kappa}(\sigma\mathbf{w},\sigma\mathbf{v})\allowbreak={\kappa}(\mathbf{w},\mathbf{v})$ for all $\mathbf{w},\mathbf{v}\in\mathbb{R}^{d},\sigma\in S_{d}$, with $S_{d}$ acting on $\mathbb{R}^{d}$ by $(\sigma\mathbf{w})_{i}=\mathbf{w}_{\sigma^{-1}(i)}$.
Since the Euclidean inner product $\langle\cdot,\cdot\rangle$ is $O(d)$-invariant and since the action of $S_{d}$ used in Definition 1 is orthogonal, ${\kappa}(\mathbf{w},\mathbf{v})=\langle\mathbf{w},\mathbf{v}\rangle^{n}$ is permutation-invariant for any $n\in\mathbb{N}$. In a similar vein, any kernel expressed in terms of $\langle\mathbf{w},\mathbf{v}\rangle$, such as ${\kappa}_{\mathcal{N}}$, is permutation-invariant.
By a direct extension of the action of $S_{d}$, we let a pair of permutations $\sigma=(\sigma_{1},\sigma_{2})\in S_{k}\times S_{d}$ act on a matrix $A=\left[A_{ij}\right]\in M(k,d)$ by
$\left(\sigma A\right)_{ij}\coloneqq A_{\sigma_{1}^{-1}(i),\sigma_{2}^{-1}(j)}.$ (4.14)
Thus, $\sigma_{1}$ acts by permuting the rows of $A$, and $\sigma_{2}$ by permuting the columns of $A$. The largest subgroup of $S_{k}\times S_{d}$ fixing $A\in M(k,d)$, namely, the isotropy subgroup of $A$, is denoted by $(S_{k}\times S_{d})_{A}$. Likewise, $(S_{d})_{\mathbf{x}}$ is the largest subgroup of $S_{d}$ fixing $\mathbf{x}\in\mathbb{R}^{d}$. For example, the isotropy subgroup of $I_{d}\in M(d,d)$ is $\Delta S_{d}$, $\Delta$ mapping a subgroup $G\subseteq S_{d}$ to the diagonal subgroup $\Delta G\coloneqq\{(g,g):g\in G\}$. The isotropy group of $\mathcal{I}_{d,1}$, $\mathcal{I}_{d_{1},d_{2}}$ generally denoting the $d_{1}\times d_{2}$-matrix with all entries equal to 1, is $S_{d}$. We denote the projection of a subgroup $H\subseteq S_{d_{1}}\times S_{d_{2}}$ onto its $i$’th component by $\pi_{i}(H)$. In particular, $\pi_{1}(\Delta G)=\pi_{2}(\Delta G)=G$ for any subgroup $G\subseteq S_{d}$.
Proposition 3 (Notation and assumption as above).
Fix $V\in M(h,d)$, $\boldsymbol{\beta}\in\mathbb{R}^{h}$, $\boldsymbol{\alpha}\in\mathbb{R}^{k}$ and define $\Gamma_{\boldsymbol{\beta},V}\coloneqq\Pi_{2}((S_{h}\times S_{d})_{V}\cap((S_{h})_{\boldsymbol{\beta}}\times S_{d}))$. If a kernel $\kappa$} is permutation-invariant then ${\mathcal{L}}_{\kappa}$ is $(S_{k})_{\boldsymbol{\alpha}}\times\Gamma_{\boldsymbol{\beta},V}$-invariant with respect to $W$. That is, ${\mathcal{L}}_{\kappa}(\sigma W,\boldsymbol{\alpha};V,\boldsymbol{\beta})={\mathcal{L}}_{\kappa}(W,\boldsymbol{\alpha};V,\boldsymbol{\beta})$ for all $\sigma\in(S_{k})_{\boldsymbol{\alpha}}\times\Gamma_{\boldsymbol{\beta},V}$.
Proposition 3 generalizes the invariance results given in [14, 36], giving an explicit description of the invariance properties of permutation-invariant kernels in terms of the symmetry of $V$, $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$. It is seen that the larger the isotropy of the target matrix $V$, the richer the invariance properties of ${\mathcal{L}}_{\kappa}$. For example, if $\boldsymbol{\alpha}=\boldsymbol{\beta}=\mathcal{I}_{d,1}$ and $V\in M(d,d)$ is a circulant matrix then $(S_{d}\times S_{d})_{V}\supseteq\Delta\mathbb{Z}_{d}$, with $\mathbb{Z}_{d}\coloneqq\mathbb{Z}/d\mathbb{Z}$ denoting the group of all cyclic permutations, rendering ${\mathcal{L}}_{\kappa}$ $S_{k}\times\mathbb{Z}_{d}$-invariant. A similar argument gives the following.
Corollary 1 (Notation and assumptions as above).
If $V=I_{d}\in M(d,d)$, $\boldsymbol{\beta}=\mathcal{I}_{d,1}$ (hence defining $T_{\mathbf{e}}$) and $\boldsymbol{\alpha}=\mathcal{I}_{k,1}$, then optimization problem (1.1) is $S_{k}\times S_{d}$-invariant under the Frobenius as well as the cubic-Gaussian tensor norms.
The corollary follows by a direct application of Proposition 3, noting that $\Gamma_{\boldsymbol{\beta},V}=\pi_{2}((S_{d}\times S_{d})_{V}\cap((S_{d})_{\boldsymbol{\beta}}\times S_{d}))=\pi_{2}(\Delta S_{d}\cap(S_{d}\times S_{d}))=\pi_{2}(\Delta S_{d})=S_{d}$.
Given $G\subseteq S_{k}\times S_{d}$, the gradient vector field of a $G$-invariant differentiable function is $G$-equivariant (an inner product on $M(k,d)$ is tacitly understood) and so maps $M(k,d)^{G}$ to itself. By Proposition 3, ${\mathcal{L}}_{\kappa}$ is $(S_{k})_{\boldsymbol{\alpha}}\times\Gamma_{\boldsymbol{\beta},V}$-invariant with respect to $W$. In particular, $\nabla_{W}{\mathcal{L}}_{\kappa}$ restricted to $M\coloneqq M(k,d)^{(S_{k})_{\boldsymbol{\alpha}}\times\Gamma_{\boldsymbol{\beta},V}}$ satisfies
$(\nabla_{W}{\mathcal{L}}_{\kappa})|_{M}(\sigma W,\boldsymbol{\alpha};V,\boldsymbol{\beta})=\sigma(\nabla_{W}{\mathcal{L}}_{\kappa})|_{M}(W,\boldsymbol{\alpha};V,\boldsymbol{\beta}),$ (4.15)
for all $\sigma\in(S_{k})_{\boldsymbol{\alpha}}\times\Gamma_{\boldsymbol{\beta},V}$.
5. Families of symmetry breaking critical points
By Corollary 1, if $k=d$ and if the target tensor used is $T_{\mathbf{e}}$ then $\mathcal{L}$} is $S_{d}\times S_{d}$-invariant. As indicated earlier, the isotropy groups of critical points of $\mathcal{L}$} found by numerical methods are typically proper subgroups of $S_{d}\times S_{d}$ and so are symmetry breaking [14]. Below, we outline a framework developed in [38, 27, 25, 36, 20] for studying phenomena of symmetry breaking of this nature. Our presentation makes extensive use of a Newton polyhedron argument which is new to the framework and allows a systematic investigation of all critical points occurring in a given fixed points space. The methods are first illustrated with reference to the $S_{d}\times S_{d}$-isotropy case under the Frobenius norm, and are later described for general isotropy groups and tensor norms. Throughout, we assume that the target tensor is $T_{\mathbf{e}}$ (but see Section A).
Fix $d\in\mathbb{N}$ and let $\mathfrak{C}(d)\in M(d,d)^{S_{d}\times S_{d}}$ be a critical point of ${\mathcal{L}}_{F}$. Since $M(d,d)^{S_{d}\times S_{d}}=\{\xi\mathcal{I}_{d,d}~|~\xi\in\mathbb{R}\}$, $\mathfrak{C}(d)=\xi(d)\mathcal{I}_{d,d}$ for some real scalar $\xi(d)\in\mathbb{R}$. The condition $\nabla{\mathcal{L}}_{F}(\mathfrak{C}(d))=0$, omitting the dependence of ${\mathcal{L}}_{F}$ on $V,\boldsymbol{\alpha}$ and $\boldsymbol{\beta}$ for convenience, reads
$\nabla{\mathcal{L}}_{F}(\xi(d)\mathcal{I}_{d,d})=0,$
giving a single polynomial equation in $\xi$ and $d$,
$\xi^{5}d^{3}-\xi^{2}=0.$ (5.16)
Thus, for any fixed $d$, there are two (real) critical points with isotropy $S_{d}\times S_{d}$: $\mathfrak{C}_{0}(d)=0_{d,d}$ and $\mathfrak{C}_{1}(d)=d^{-1}\mathcal{I}_{d,d}$. Moreover,
${\mathcal{L}}_{F}(\xi\mathcal{I}_{d,d})$ $=\xi^{6}d^{5}-2\xi^{3}d^{2}+d,$ (5.17)
and so the respective loss is ${\mathcal{L}}_{F}(\mathfrak{C}_{0}(d))=d$ and ${\mathcal{L}}_{F}(\mathfrak{C}_{1}(d))=d-\frac{1}{d}$. The loss, in both cases, is strictly positive for $d\geq 2$ and so in particular $\mathfrak{C}_{1}$ cannot be a global minimum. To determine whether $\mathfrak{C}_{1}$ defines a local minimum, a local maximum or a saddle, we compute the Hessian spectrum. The symmetry of $\mathfrak{C}_{1}$ allows a considerable simplification in the computation of the Hessian spectrum using results from the representation theory of the symmetric group [39]. The analysis proceeds by considering four irreducible $S_{d}$-representations (isomorphism classes of $S_{d}$-invariant linear spaces containing no proper $S_{d}$-invariant subspace but the trivial):
  • The trivial representation $\mathfrak{t}_{d}$ of degree 1.
  • The standard representation $\mathfrak{s}_{d}$ of $S_{d}$ of degree d1d-1.
  • The exterior square representation $\mathfrak{x}_{d}=\wedge^{2}\mathfrak{s}_{d}$ of degree $\frac{(d-1)(d-2)}{2}$.
  • The representation $\mathfrak{y}_{d}$ associated to the partition (d2,2)(d-2,2) of degree $\frac{d(d-3)}{2}$.
The Hessian spectrum is computed and given separately for each irreducible representation or tensor products thereof (see Table 1 and Table 3). We refer to [17, Section XII] for more details on the approach.
In the general case, one considers sequences of subgroups $(G_{d})_{d\in\mathbb{N}}$ of $S_{k}\times S_{d}$ for which $\dim(M(k,d)^{G_{d}})$ stabilizes for sufficiently large $d$. To simplify exposition, assume $\dim(M(k,d))$ is fixed and equal to $N\in\mathbb{N}$ for all $d\in\mathbb{N}$. The fixed point space bears a natural parameterization given by a linear isomorphism $\Xi\coloneqq\Xi^{G_{d}}:\mathbb{R}^{N}\to M(k,d)^{G_{d}}$. Families of critical points in $M(k,d)^{G_{d}}$ may therefore be expressed by $\mathfrak{C}(d)=\Xi(\mathfrak{c}(d))$, with $\mathfrak{c}(d)\in\mathbb{R}^{N}$ satisfying a system of $N$ polynomial equations,
$\Xi^{-1}\circ\nabla{\mathcal{L}}|_{M(k,d)^{G_{d}}}\circ\Xi(\mathfrak{c}(d))=0.$ (5.18)
Note that pulling $\nabla{\mathcal{L}}|_{M(k,d)^{G_{d}}}$ back to $\mathbb{R}^{N}$ via $\Xi^{-1}$ is possible by the $G_{d}$-equivariance of the gradient vector field $\nabla{\mathcal{L}}$, see (4.15). The single polynomial equation given in the $S_{d}\times S_{d}$-case, namely Equation 5.16, now becomes a system of polynomial equations, and so expressing $\mathfrak{C}(d)$ in terms of simple formulae in $d$ is generally hard. However, for the purpose of characterizing the stability of families of critical points it suffices, in many cases, to estimate the Hessian spectrum to a certain order. An exact computation of $\mathfrak{C}(d)$ is not required—estimates to a limited order suffice. Here estimates are obtained by considering partial sums of Puiseux series constructed by repeating a Newton polytope argument (details appear in Section 5.1). In Section C, we provide a detailed construction of a family of (complex) critical points in Puiseux series in $d^{\frac{-1}{4}}$ based on a partial (fractional) sum so obtained. General results on the existence of Puiseux series solutions are given in [40]. To our applications, the CSL, applied to the semialgebraic solution set (5.18), suffices [15]. A more direct approach is given by applying the CSL to the solution set transformed by (e.g., [41])
${\boldsymbol{\xi}}\coloneqq(\xi_{1},\dots,\xi_{N})$ $\mapsto\left(\frac{\xi_{1}}{\sqrt{1+\|{\boldsymbol{\xi}}\|^{2}}},\dots,\frac{\xi_{N}}{\sqrt{1+\|{\boldsymbol{\xi}}\|^{2}}},\frac{1}{\sqrt{1+\|{\boldsymbol{\xi}}\|^{2}}}\right),$ (5.19)
see [15] (also addressing the dimensionality of strata of the solution set).
5.1. Critical points under the Frobenius norm
For the target tensor $T_{\mathbf{e}}$, the loss function ${\mathcal{L}}_{F}$ takes the form
$\varphi(W)\coloneqq\left\|\sum_{i=1}\mathbf{w}_{i}^{\otimes{}3}-\sum_{i=1}\mathbf{e}_{i}^{\otimes{}3}\right\|_{F}^{2}.$ (5.20)
Isotropy $\Delta S_{d}$.
Critical points of $S_{d}\times S_{d}$-isotropy are given in the previous section and so we begin by considering the $\Delta S_{d}$-case. The fixed point space corresponding to $\Delta S_{d}$ is
$M(d,d)^{\Delta S_{d}}\coloneqq\{\xi_{1}I_{d}+\xi_{2}(\mathcal{I}_{d,d}-I_{d})~|~\xi_{i}\in\mathbb{R}\}.$
By the preceding section, we seek a solution of the form
$\xi_{i}(d)=A_{i}d^{\tau_{i}}+\text{terms with lower exponents},i=1,2,$ (5.21)
with $A_{i}\neq 0$ (vanishing $A_{i}$ will be discussed later) and $\tau_{i}\in\mathbb{Q}$ (note that exponents are assumed decreasing) for a given family of critical points. The critical point equations corresponding to the restriction of $\nabla{\varphi}$} to $M(d,d)^{\Delta S_{d}}$ (divided by 2) are
$0$ $=3\xi_{2}^{5}d^{3}+d^{2}\left(15\xi_{1}\xi_{2}^{4}-15\xi_{2}^{5}\right)+d\left(6\xi_{1}^{3}\xi_{2}^{2}+12\xi_{1}^{2}\xi_{2}^{3}-42\xi_{1}\xi_{2}^{4}+24\xi_{2}^{5}\right)$
$+3\xi_{1}^{5}-6\xi_{1}^{3}\xi_{2}^{2}-12\xi_{1}^{2}\xi_{2}^{3}-3\xi_{1}^{2}+27\xi_{1}\xi_{2}^{4}-12\xi_{2}^{5},$
$0$ $=3\xi_{2}^{5}d^{3}+d^{2}\left(15\xi_{1}\xi_{2}^{4}-15\xi_{2}^{5}\right)+d\left(30\xi_{1}^{2}\xi_{2}^{3}-60\xi_{1}\xi_{2}^{4}+30\xi_{2}^{5}\right)$
$+3\xi_{1}^{4}\xi_{2}+12\xi_{1}^{3}\xi_{2}^{2}-54\xi_{1}^{2}\xi_{2}^{3}+60\xi_{1}\xi_{2}^{4}-21\xi_{2}^{5}-3\xi_{2}^{2}.$ (5.22)
When $\xi_{i}$ in the above are substituted by (5.21), every term of maximal total degree must appear at least twice so that a cancellation by a different term of the same (total) degree is possible by a suitable choice of $A_{i}$, see e.g., [42]. Equivalently, arguments maximizing
$\max\left(2\tau_{1},5\tau_{1},5\tau_{2}+3,\tau_{1}+4\tau_{2}+2,2\tau_{1}+3\tau_{2}+1,3\tau_{1}+2\tau_{2}+1\right)$ (5.23)
must be attained at least twice, as well as arguments maximizing
$\max\left(2\tau_{2},3\tau_{1}+2\tau_{2},4\tau_{1}+\tau_{2},5\tau_{2}+3,\tau_{1}+4\tau_{2}+2,2\tau_{1}+3\tau_{2}+1\right).$
The only possibilities are $(\tau_{1},\tau_{2})=(0,-3/4)$ and $(\tau_{1},\tau_{2})=(-1,-1)$. Since $A_{i}$ are chosen so that terms of maximal total degree cancel out, for $(\tau_{1},\tau_{2})=(-1,-1)$, $(A_{1},A_{2})=(\pm 1,1)$, defining two families of minima: $\mathfrak{C}_{1}$ of isotropy $S_{d}\times S_{d}$ encountered earlier and $\mathfrak{C}_{2}(d)=(1/d)(\mathcal{I}_{d,d}-2I_{d})+o(d)$ of isotropy $\Delta S_{d}$. For $(\tau_{1},\tau_{2})=(0,-3/4)$, $(A_{1},A_{2})=(1,\sqrt[4]{-1})$, where $\sqrt[4]{-1}$ denotes any fourth root of 1-1, none of which is real. The families of critical points $\mathfrak{C}_{0}(d)\coloneqq 0_{d,d}$ and $\mathfrak{C}_{I}(d)\coloneqq I_{d,d}$ are obtained by letting either $A_{i}$ or both vanish. Additional terms of the Puiseux series are computed by repeating the process.
Remark 4.
Depending on the application, tensor decomposition problems may be studied over the complex numbers, e.g., [43, 44, 45, 46]. The symmetry breaking framework carries over to this setting without change. For example, a family of complex $\Delta S_{d}$-critical points corresponding to the pair of exponents $(\tau_{1},\tau_{2})=(0,-3/4)$ is given by
$(\mathfrak{c}_{\mathbb{C}})_{1}(d)$ $=1-\frac{2i}{3d^{\frac{1}{2}}}+\frac{\sqrt{2}\left(1+i\right)}{d^{\frac{3}{4}}}-\frac{77}{36d}+o(d),$
$(\mathfrak{c}_{\mathbb{C}})_{2}(d)$ $=\frac{\sqrt{2}\left(1+i\right)}{2d^{\frac{3}{4}}}-\frac{5}{4d}+o(d).$
Note that here Puiseux series are given in terms of $d^{-\frac{1}{4}}$.
Remark 5.
By Bézout’s theorem the number of solutions of (5.1) is generically 25 (the product of the total degree of the equations), exceeding the actual number of solutions. The more refined bound given by Bernstein’s theorem [47] is still lose. In both cases, this results from additional structure on account of symmetry, see [15].
Isotropy groups with two or more factors.
For critical points of isotropy $\Delta(S_{d-1}\times S_{1})$, the analog for (5.23) comprises five piecewise-linear conditions (with seventy two terms). A $\Delta(S_{d-1}\times S_{1})$-family of critical points, denoted $\mathfrak{C}_{3}=\Xi^{\Delta(S_{d-1}\times S_{1})}(\mathfrak{c}_{3})$, is given by
$(\mathfrak{c}_{3})_{1}(d)$ $=-\frac{1}{d}-\frac{13}{3d^{2}}-\frac{77}{9d^{3}}+\frac{6421}{81d^{4}}+o(d^{-4}),$ $(\mathfrak{c}_{3})_{4}(d)$ $=o(d^{-4}),$
$(\mathfrak{c}_{3})_{2}(d)$ $=\frac{1}{d}+\frac{13}{3d^{2}}+\frac{149}{9d^{3}}+\frac{2867}{81d^{4}}+o(d^{-4}),$ $(\mathfrak{c}_{3})_{5}(d)$ $=1+o(d^{-4}),$
$(\mathfrak{c}_{3})_{3}(d)$ $=o(d^{-4}).$
In some cases, the entries of a family of critical points do not depend on $d$. One such example is $\mathfrak{C}_{4}=\Xi^{\Delta(S_{d-2}\times S_{2})}(\mathfrak{c}_{4})$ where
$(\mathfrak{c}_{4})_{1}(d)=1,~(\mathfrak{c}_{4})_{5}(d)=1/2,~(\mathfrak{c}_{4})_{6}(d)=1/2$ (5.24)
(only non-zero entries are described). Additional examples are given by $\mathfrak{C}_{0}$, $\mathfrak{C}_{I}$ and $\mathfrak{C}_{5,t}$, defined in Section 2.1.
5.2. Critical points for the Cubic-Gaussian norm
As earlier, we assume that the target tensor is $T_{\mathbf{e}}$. The loss function ${\mathcal{L}}_{\mathcal{N}}$ takes the form
$\psi(W)\coloneqq\left\|\sum_{i=1}\mathbf{w}_{i}^{\otimes{}3}-\sum_{i=1}\mathbf{e}_{i}^{\otimes{}3}\right\|_{\mathcal{N}}^{2}.$ (5.25)
Isotropy $S_{d}\times S_{d}$.
We begin by considering critical points in $M(d,d)^{S_{d}\times S_{d}}=\{\xi_{1}\mathcal{I}_{d,d}~|~\xi\in\mathbb{R}\}$. We have,
$\psi(\xi I)$ $=15\xi^{6}d^{5}-18\xi^{3}\left(d^{3}+\frac{2d^{2}}{3}\right)+15d,$
$(\nabla\psi)|_{M(d,d)^{S_{d}\times S_{d}}}$ $=90\xi^{5}d^{3}-54\xi^{2}\left(d+\frac{2}{3}\right).$
There are therefore two families of critical points corresponding to isotropy $S_{d}\times S_{d}$, $\mathfrak{D}_{0}(d)=0_{d,d}$ and
$\mathfrak{D}_{1}(d)=\sqrt[3]{\frac{3}{5d^{3}}\left(d+\frac{2}{3}\right)}\cdot\mathcal{I}_{d,d}.$
The loss at $\mathfrak{D}_{0}$ is $\varphi(\mathfrak{D}_{0}(d))=15d$. The Hessian at $\mathfrak{D}_{0}$ vanishes (but similarly to $\mathfrak{C}_{0}$, the tensor of third-order partial derivatives does not vanish). The loss at $\mathcal{D}_{1}$ is $\psi(\mathcal{D}_{1})=\frac{48d}{5}-\frac{36}{5}-\frac{12}{5d}$. The Hessian spectrum is given in Table 3.
Diagonal Isotropy groups.
For isotropy $\Delta S_{d}$, the system of piecewise-linear conditions analogous to (5.23) implies that a critical point given by (5.21) must have $(\tau_{1},\tau_{2})=(0,-1)$, $(\tau_{1},\tau_{2})=(-1/6,-2/3)$ or $(\tau_{1},\tau_{2})=(-2/3,-2/3)$. The first and the second case imply $A_{1}A_{2}=0$, and are therefore disregarded (recall that $A_{i}$ are assumed to be non-zero). The third case gives $\mathfrak{D}_{1}$.
The family of critical points for which $A_{2}=0$ is $\mathfrak{D}_{I}(d)=I_{d}$. Clearly, $\psi(\mathfrak{D}_{I})=0$. The Hessian spectrum is given in Table 3. A family of $\Delta(S_{d-1}\times S_{1})$-critical points $\mathfrak{D}_{2}=\Xi^{\Delta(S_{d-1}\times S_{1})}(\mathfrak{d}_{2})$ giving spurious minima is defined, modulo $o(d^{-4})$-terms, by
$(\mathfrak{d}_{2})_{1}(d)$ $=1-\frac{5}{3d^{3}}+\frac{9}{d^{4}},$ $(\mathfrak{d}_{2})_{4}(d)=0,$
$(\mathfrak{d}_{2})_{2}(d)$ $=-\frac{1}{d^{3}}+\frac{7}{d^{4}},$ $(\mathfrak{d}_{2})_{5}(d)=0,$
$(\mathfrak{d}_{2})_{3}(d)$ $=\frac{1}{d}-\frac{1}{d^{2}}+\frac{2}{d^{3}}+\frac{10}{3d^{4}}.$
6. Concluding remarks
Standard approaches [6, 7, 8, 9, 10, 11] for decomposing real symmetric tensors often fail to return a valid decomposition [12, 10, 11, 13] of nonconvexity of the associated optimization landscape. Our focus in this article has been on studying in some depth obstructions to local optimization methods using techniques pertaining to symmetry breaking [15]. The complex array of saddles and minima so revealed exhibits various phenomena that influence the efficacy of local optimization methods. Some phenomena are desirable; for instance, saddles with higher objective values tending to exhibit a greater number of descent directions, as well as the emergence of such directions under over-parametrization (see op. cit. and [26]). Others are awkward, as third-order saddles existing in the optimization landscape, which can unduly slow down the optimization process or even cause an optimization method to erroneously declare convergence to a local minimum. Certain behaviors are common to different tensor norms, such as highly skewed Hessian spectra, while others, like the growth rates of Hessian eigenvalues, differ. These phenomena, along with others discussed in detail in Section 2, help clarify key questions foundational to theories that aim to characterize the decomposability of symmetric tensors using local optimization methods. We conclude by noting that, although our emphasis has been on orthonormally decomposable tensors, this assumption is not strictly necessary. Variations, beyond those discussed in Section 2, include, for instance, further emphasis on collinearity factors [48], over-factoring [13], and tensors with generic rank [49].
In a broader context, the applicability of the symmetry-breaking framework stems from the presence of critical points whose isotropy groups are large. A formal mathematical mechanism accounting for the emergence of such symmetry-breaking points is developed in [15].
Acknowledgements
We thank Michael Field, Yuval Peled and Avi Wigderson for several helpful and insightful discussions. The research was supported by the Israel Science Foundation (grant No. 724/22).
Appendix A Generalities
Following Section 4, we consider a formulation of the objective function ${\mathcal{L}}_{\kappa}:M(k,d)\to\mathbb{R}$ (having dependence on other problem parameters implicit) given by a kernel ${\kappa}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ satisfying ${\kappa}(\mathbf{w},\mathbf{v})={\kappa}(\mathbf{v},\mathbf{w})$,
${\mathcal{L}}$ $(W)=\sum_{i,j=1}^{k}{\kappa}(\mathbf{w}_{i},\mathbf{w}_{j})-2\sum_{i=1}^{k}\sum_{j=1}^{h}{\kappa}(\mathbf{w}_{i},\mathbf{v}_{j})+\sum_{i,j=1}^{h}{\kappa}(\mathbf{v}_{i},\mathbf{v}_{j}),$ (A.26)
where $\mathbf{w}_{i}$ (resp. $\mathbf{v}_{i}$) is a column vector denoting the $i$’th row of $W\in M(k,d)$ (resp. $V\in M(h,d)$). The relevant literature considers different conventions for arranging first-, second- and higher-order derivatives of a real-valued function defined over matrices, see e.g., [50]. Here we find it convenient to identify $M(k,d)$ with $\mathbb{R}^{kd}$ and regard $W\in M(k,d)$ as a column vector having $kd$ entries given by the rows of $W$ transposed and stacked on top of one another. Thus, the differential $D{\mathcal{L}}_{\kappa}$ maps $\mathbb{R}^{kd}$ to $L(\mathbb{R}^{kd},\mathbb{R})$, where $L(E,F)$ generally denotes the space of all bounded linear operators from a vector space $E$ to a vector space $F$. The gradient $\nabla{\mathcal{L}}_{\kappa}:\mathbb{R}^{kd}\to\mathbb{R}^{kd}$ (defined by the usual identification) is given by
$\nabla{\mathcal{L}}(W)$ $=2\sum_{i=1}^{k}\mathbf{e}_{i}\otimes\left(\sum_{j=1}^{k}\kappa_{\mathbf{w}}(\mathbf{w}_{i},\mathbf{w}_{j})-\sum_{j=1}^{h}\kappa_{\mathbf{w}}(\mathbf{w}_{i},\mathbf{v}_{j})\right),$ (A.27)
where $\mathbf{e}_{i}$ denotes the $i$-th unit vector and $\kappa_{\mathbf{w}}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ a column vector containing the partial derivatives of $\kappa$ with respect to the entries of the first argument. The Hessian $\nabla{\mathcal{L}}^{2}_{\kappa}:\mathbb{R}^{kd}\to L(\mathbb{R}^{kd},L(\mathbb{R}^{kd},\mathbb{R}))\cong M(\mathbb{R}^{kd},\mathbb{R}^{kd})$ is given by
$\nabla^{2}{\mathcal{L}}(W)$ $=2\sum_{i,j=1}^{k}\mathbf{e}_{i}\mathbf{e}_{j}^{\top}\otimes\kappa_{\mathbf{w},\mathbf{v}}(\mathbf{w}_{i},\mathbf{w}_{j})$ (A.28)
$+2\sum_{i=1}^{k}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\otimes\left(\sum_{j=1}^{k}\kappa_{\mathbf{w},\mathbf{w}}(\mathbf{w}_{i},\mathbf{w}_{j})-\sum_{j=1}^{h}\kappa_{\mathbf{w},\mathbf{w}}(\mathbf{w}_{i},\mathbf{v}_{j})\right),$
where $\kappa_{\mathbf{w},\mathbf{w}}$ (resp. $\kappa_{\mathbf{w},\mathbf{v}}$) is the $d\times d$ Jacobian matrix of $\kappa_{\mathbf{w}}$ with respect to $\mathbf{w}$ (resp. $\mathbf{v}$).
If ${\kappa}(\mathbf{w},\mathbf{v})=\langle\mathbf{w},\mathbf{v}\rangle^{n}$ and $h=d$, Equation (A.26) reads
${\mathcal{L}}$ $(W)=\sum_{i,j=1}^{k}\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{n}-2\sum_{i=1}^{k}\sum_{j=1}^{d}\langle\mathbf{w}_{i},\mathbf{v}_{j}\rangle^{n}+\sum_{i,j=1}^{d}\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle^{n}.$ (A.29)
If $\mathbf{v}_{1},\dots,\mathbf{v}_{d}$ are orthonormal, $V$ denotes the matrix whose rows are $\mathbf{v}_{i}$ and $\hat{\mathbf{w}}_{i}\coloneqq({\mathbf{w}}_{i}V)^{\top}$, then
$\hat{{\mathcal{L}}}$ $(W)\coloneqq{\mathcal{L}}(WV)=\sum_{i,j=1}^{k}\langle\hat{\mathbf{w}}_{i},\hat{\mathbf{w}}_{j}\rangle^{n}-2\sum_{i=1}^{k}\sum_{j=1}^{d}\langle\hat{\mathbf{w}}_{i},\mathbf{e}_{j}\rangle^{n}+\sum_{i,j=1}^{d}\langle\mathbf{e}_{i},\mathbf{e}_{j}\rangle^{n}.$ (A.30)
Thus, the optimization landscape of $\mathcal{L}$}} and $\hat{{\mathcal{L}}}$, the latter corresponding to the target tensor $T_{\mathbf{e}}$, are identical modulo orthonormal transformation of the domain (a similar argument shows that this applies to Gauss-cubic norm as well). It is therefore no loss of generality to focus on the optimization landscape of $\hat{{\mathcal{L}}}$. Equation (A.27) now reads
$\nabla{\mathcal{L}}(W)$ $=2n\sum_{i=1}^{k}\mathbf{e}_{i}\otimes\left(\sum_{j=1}^{k}\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{n-1}\mathbf{w}_{j}-\sum_{j=1}^{d}\langle\mathbf{w}_{i},\mathbf{e}_{j}\rangle^{n-1}\mathbf{e}_{j}\right).$ (A.31)
It follows by the above shows that eigenvectors of the target tensor $T$ [51, 52, 16] give critical points. Indeed, a vector $\mathbf{w}\in\mathbb{R}^{d}$ is an eigenvector of $T$ if $T\mathbf{w}^{n-1}=\lambda\mathbf{w}$ for some $\lambda\in\mathbb{R}$, where $T\mathbf{w}^{n-1}$ generally denotes a $d$-dimensional vector given by $(T\mathbf{w}^{n-1})_{i}\coloneqq\sum^{d}_{i_{2},\dots,i_{n}=1}T_{i,i_{2},\dots,i_{n}}w_{i_{2}}\cdots w_{i_{n}}$. Given an eigenpair $(\mathbf{w},\lambda)$, define a matrix $W_{t}$ whose first row is $t\mathbf{w},t\in\mathbb{R}$, and is otherwise zero. We have,
$\nabla{\mathcal{L}}(W_{t})$ $=2n\mathbf{e}_{1}\otimes\left(t^{2n-1}\langle\mathbf{w},\mathbf{w}\rangle^{n-1}\mathbf{w}-t^{n-1}\lambda\mathbf{w}\right).$ (A.32)
Thus, $W_{t}$ defines a critical point for $t=(\lambda/\langle\mathbf{w},\mathbf{w}\rangle^{n-1})^{1/n}$. Generalizing the latter argument, if $r$ orthogonal eigenvectors $\mathbf{w}_{1},\dots,\mathbf{w}_{r}$ are given, $t_{i}$ may be set individually as in (A.32). Forming a weight matrix $W_{(r)}$ whose first $r$ rows are $t_{i}\mathbf{w}_{1},\dots,t_{r}\mathbf{w}_{r}$ and is otherwise zero, we have,
$\nabla{\mathcal{L}}(W_{(r)})$ $=2n\sum_{i=1}^{r}\mathbf{e}_{i}\otimes\left(t_{i}^{2n-1}\langle\mathbf{w}_{i},\mathbf{w}_{i}\rangle^{n-1}\mathbf{w}_{i}-t_{i}^{n-1}\lambda\mathbf{w}_{i}\right)=0.$ (A.33)
Of course, families of critical points whose weight matrices are not orthonormal, such as $\mathfrak{C}_{2}$, do not arise by this mechanism.
Remark 6.
An equivalent way of defining an eigenvector follows using $f_{T}:\mathbb{R}^{d}\to\mathbb{R}:w\mapsto\langle T,w^{d}\rangle_{F}$. A simple argument shows that
$(\mathbf{w},\lambda)$ is an eigenpair iff it satisfies
$\nabla f_{T}(w)=d\lambda w.$
It is possible to replace the Frobenius inner product in definition of $f_{T}$ with a different inner product and obtain similar results.
The expressions of the Hessian (A.28) are given by,
$\nabla^{2}$ ${\mathcal{L}}(W)=2\sum_{i,j=1}^{k}\mathbf{e}_{i}\mathbf{e}_{j}^{\top}\otimes(n\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{n-1}I_{d}+n(n-1)\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{n-2}\mathbf{w}_{j}\mathbf{w}_{i}^{\top})$ (A.34)
$+2\sum_{i=1}^{k}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\otimes\left(\sum_{j=1}^{k}n(n-1)\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{n-2}\mathbf{w}_{j}\mathbf{w}_{j}^{\top}-\sum_{j=1}^{d}n(n-1)\langle\mathbf{w}_{i},\mathbf{v}_{j}\rangle^{n-2}\mathbf{v}_{j}\mathbf{v}_{j}^{\top}\right).$
Analytic expressions for the Gauss-cubic norm are derived similarly.
Appendix B Proofs
B.1. Proofs for Section 2.1
Proof of Proposition 1
For any mode of the target tensor $T_{\mathbf{e}}=\sum_{i=1}^{d}\mathbf{e}_{i}^{\otimes 3}$, the matrix containing the respective columns is simply the identity matrix $I_{d}$. The rank of the identity matrix $I_{d}$ is $d$ and so trivially every $d$ columns are independent. Thus, by Kruskal’s identifiability result [18], $T_{\mathbf{e}}$ admits a unique decomposition modulo permutation and scaling of the columns if $3d\geq 2d+2$, i.e., if $d\geq 2$.
We also provide the details for $T=\sum_{i=1}^{d}\mathbf{e}_{i}^{\otimes n},n\geq 4$. For even $n$, we assume that $\boldsymbol{\alpha}$ is adjusted during the optimization process. If $(W,\boldsymbol{\alpha})$ is a global minimizer, then ${\mathcal{L}}(W)=0$ and so independently of the underlying tensor norm, $T=\sum_{i=1}^{d}\alpha_{i}\mathbf{w}_{i}^{\otimes n}$. Rearranging indices if needed, we may assume $\mathbf{w}_{i}=c_{i}\mathbf{e}_{i},~c_{i}\in\mathbb{R}$ for $i\in[d]$, giving
$\alpha_{j}c_{j}^{n}=\left(\sum_{i=1}^{d}\alpha_{i}c_{i}^{n}\mathbf{e}_{i}^{\otimes n}\right)_{j,\dots,j}=\left(\sum_{i=1}^{d}\alpha_{i}\mathbf{w}_{i}^{\otimes n}\right)_{j,\dots,j}=T_{j,\dots,j}=1.$
Therefore, if $n$ is odd, $c_{j}=\alpha_{j}^{-1/n}$. If $n$ is even, necessarily $\alpha_{j}\geq 0$ and $c_{j}=\pm\alpha_{j}^{-1/n}$. (For $n=3$, we assumed $\alpha_{j}=1$ and so $c_{j}=1$.)
B.2. Proof of Proposition 2
The proposition follows by a direct computation. First, we show that $\mathfrak{C}_{5,t}(d)\coloneqq I_{d-2}\oplus\begin{bmatrix}(1-t^{3})^{1/3}&0\\ t&0\end{bmatrix}$ is a critical point for every $t\in\mathbb{R}$ and $d\in\mathbb{N}$. As usual, we let $\mathbf{w}_{i}$ denote the rows of $\mathfrak{C}_{5,t}$. Referring to Equation A.33, since the first d2d-2 rows are orthogonal to other rows, we have
$\nabla{\mathcal{L}}(\mathfrak{C}_{5,t}(d))$ $=2n\sum_{i=d-1}^{d}\mathbf{e}_{i}\otimes\left(\sum_{i=d-1}^{d}\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{2}\mathbf{w}_{j}-\langle\mathbf{w}_{i},\mathbf{e}_{d-1}\rangle^{2}\mathbf{e}_{d-1}\right).$
The term in the above corresponding to i=d1i=d-1 is $\left((1-t^{3})^{2/3}\right)^{2}(1-t^{3})^{1/3}+\left(t(1-t^{3})^{1/3}\right)^{2}t-(1-t^{3})^{2/3}=0$. Similarly, for $i=d$, $t^{5}+\left((1-t^{3})^{1/3}t\right)^{2}(1-t^{3})^{1/3}-t^{2}=0$. Moreover, by Equation A.30,
${\mathcal{L}}(\mathfrak{C}_{5,t}(d))$ $=\sum_{i,j=1}^{d}\langle\mathbf{w}_{i},\mathbf{w}_{j}\rangle^{3}-2\sum_{i=1}^{d}\sum_{j=1}^{d}\langle\mathbf{w}_{i},\mathbf{e}_{j}\rangle^{3}+\sum_{i,j=1}^{d}\langle\mathbf{e}_{i},\mathbf{e}_{j}\rangle^{3}$
$=2t^{3}(1-t^{3})+t^{6}+1-2t^{3}+t^{6}=1.$
Therefore, the level set ${\mathcal{L}}^{-1}(1)$ is unbounded, hence $\mathcal{L}$} is not proper.
B.3. Proof of Proposition 3
Let $\sigma=(\sigma_{1},\sigma_{2})\in(S_{k})_{\boldsymbol{\alpha}}\times\Gamma_{\boldsymbol{\beta},V}=(S_{k})_{\boldsymbol{\alpha}}\times\Pi_{2}((S_{h}\times S_{d})_{V}\cap((S_{h})_{\boldsymbol{\beta}}\times S_{d}))$. Since $\sigma_{2}\in\Gamma_{\boldsymbol{\beta},V}$, there exists $\rho\in(S_{h})_{\boldsymbol{\beta}}$ such that $(\rho,\sigma_{2})\in(S_{h}\times S_{d})_{V}$. In particular, $\rho^{-1}\in(S_{h})_{\boldsymbol{\beta}}$ and $(\rho^{-1},\sigma_{2}^{-1})\in(S_{h}\times S_{d})_{V}$. By (4.11) and the permutation invariance of $\kappa$},
${\mathcal{L}}_{\kappa}$ $(\sigma W;V)$
$=\sum_{i,j=1}^{k}\alpha_{i}\alpha_{j}{\kappa}(\sigma_{2}\mathbf{w}_{\sigma_{1}^{-1}(i)},\sigma_{2}\mathbf{w}_{\sigma_{1}^{-1}(j)})-2\sum_{i=1}^{k}\sum_{j=1}^{h}\alpha_{i}\beta_{j}{\kappa}(\sigma_{2}\mathbf{w}_{\sigma_{1}^{-1}(i)},\mathbf{v}_{j})+\sum_{i,j=1}^{h}\beta_{i}\beta_{j}{\kappa}(\mathbf{v}_{i},\mathbf{v}_{j})$
$=\sum_{i,j=1}^{k}\alpha_{\sigma_{1}(i)}\alpha_{\sigma_{1}(j)}{\kappa}(\mathbf{w}_{i},\mathbf{w}_{j})-2\sum_{i=1}^{k}\sum_{j=1}^{h}\alpha_{\sigma_{1}(i)}\beta_{j}{\kappa}(\mathbf{w}_{i},\sigma_{2}^{-1}\mathbf{v}_{j})+\sum_{i,j=1}^{h}\beta_{i}\beta_{j}{\kappa}(\mathbf{v}_{i},\mathbf{v}_{j})$
$=\sum_{i,j=1}^{k}\alpha_{i}\alpha_{j}{\kappa}(\mathbf{w}_{i},\mathbf{w}_{j})-2\sum_{i=1}^{k}\sum_{j=1}^{h}\alpha_{i}\beta_{\rho^{-1}(j)}{\kappa}(\mathbf{w}_{i},\sigma_{2}^{-1}\mathbf{v}_{j})+\sum_{i,j=1}^{h}\beta_{\rho^{-1}(i)}\beta_{\rho^{-1}(j)}{\kappa}(\mathbf{v}_{i},\mathbf{v}_{j})$
$=\sum_{i,j=1}^{k}\alpha_{i}\alpha_{j}{\kappa}(\mathbf{w}_{i},\mathbf{w}_{j})-2\sum_{i=1}^{k}\sum_{j=1}^{h}\alpha_{i}\beta_{j}{\kappa}(\mathbf{w}_{i},\sigma_{2}^{-1}\mathbf{v}_{\rho(j)})+\sum_{i,j=1}^{h}\beta_{i}\beta_{j}{\kappa}(\sigma_{2}^{-1}\mathbf{v}_{\rho(i)},\sigma_{2}^{-1}\mathbf{v}_{\rho(j)})$
$={\mathcal{L}}(W;(\rho^{-1},\sigma_{2}^{-1})V)$
$={\mathcal{L}}(W;V),$
concluding the proof.
Appendix C Existence and uniqueness of Puiseux series
We describe a construction of Puiseux series using the (real or complex) analytic implicit function theorem [26, 20], see generally[53, 54, 40] for additional approaches. The method is illustrated with reference to $\mathfrak{c}_{\mathbb{C}}$ (see Remark 4) with isotropy group $\Delta S_{d}$ and so given by two Puiseux series $\xi_{1}(d)$ and $\xi_{2}(d)$. The system of critical equations restricted to the fixed point space $M(d,d)^{\Delta S_{d}}$ is given in (5.1); denote it by $E(\xi_{1},\xi_{2},d)$. Repeatedly applying a Newton polyhedron argument (see Section 5.1) gives
$\xi_{1}(d)$ $=1-\frac{2i}{3d^{\frac{2}{4}}}+\frac{\sqrt{2}\cdot\left(1+i\right)}{d^{\frac{3}{4}}}-\frac{77}{36d}+\frac{43\sqrt{2}\cdot\left(1-i\right)}{96d^{\frac{5}{4}}}+\frac{2821i}{5184d^{\frac{6}{4}}}$
$+\frac{8491\sqrt{2}\left(-1-i\right)}{18432d^{\frac{7}{4}}}+O(d^{-2}),$
$\xi_{2}(d)$ $=\frac{\sqrt{2}\cdot\left(1+i\right)}{2d^{\frac{3}{4}}}-\frac{5}{4d}+\frac{49\sqrt{2}\cdot\left(1-i\right)}{192d^{\frac{5}{4}}}+\frac{119i}{96d^{\frac{6}{4}}}+\frac{39799\sqrt{2}\left(-1-i\right)}{36864d^{\frac{7}{4}}}$
$+\frac{13}{9d^{2}}+O(d^{-\frac{9}{4}}).$
The trimmed Puiseux series are defined by
$\xi_{1}^{(n_{1})}(d)$ $\coloneqq\sum_{i=0}^{n_{1}}[d^{\frac{-i}{4}}]\xi_{1}(d),$
$\xi_{2}^{(n_{2})}(d)$ $\coloneqq\sum_{i=0}^{n_{2}}[d^{\frac{-i}{4}}]\xi_{2}(d),$
where $d^{\alpha}]P$ denotes the $\alpha$-order term of a Puiseux series $P$. Thus, for example, $\xi_{1}^{(2)}(d)=1-\frac{2i}{3d^{\frac{2}{4}}}$ and $\xi_{2}^{(2)}(d)=0$. Next, set $s=d^{\frac{-1}{4}}$. This substitution turns $\xi_{i}(d)$ into a power series having integral exponents. The set of equations so obtained (divided by 2) is
$\overline{E}(x,y,s)\coloneqq E(\xi_{1}^{(n_{1})}(s)+s^{n_{1}+1}x,\xi_{2}^{(n_{2})}(s)+s^{n_{2}+1}y,s^{-4}),$
where the trimming parameters, $n_{1}$ and $n_{2}$, are chosen so that the Jacobian of $\overline{E}$,
$\begin{pmatrix}\frac{\partial\overline{E}_{1}}{\partial x},\frac{\partial\overline{E}_{1}}{\partial y}\\ \frac{\partial\overline{E}_{2}}{\partial x},\frac{\partial\overline{E}_{2}}{\partial y}\end{pmatrix}$,
at the initial conditions $x=[s^{n_{1}+1}]\xi_{1}(s)$, $y=[s^{n_{2}+1}]\xi_{2}(s)$ and $s=0$ is non-singular (assuming momentarily that powers of $s$ occurring in each equation have been canceled). We may now invoke the analytic implicit function theorem (here, in the complex category) and so present $x(s)$ and $y(s)$ as power series in $s$ in a neighborhood of $s=0$. Reversing the substitution, that is, substituting $d^{\frac{-1}{4}}$ for $s$, completes the derivation (note, convergence is now in a neighborhood of $d=\infty$, that is, for sufficiently large $d$). For $\mathfrak{c}_{\mathbb{C}}$, suffices it to take $n_{1}=1$ and $n_{2}=2$. Thus, $\overline{E}(x,y,s)=E(1+s^{2}x,s^{3}y,s^{-4})$. We have, for example,
$\overline{E}(x,y,s)_{1}$ $=-12s^{15}y^{5}+27s^{14}xy^{4}-12s^{13}x^{2}y^{3}+s^{12}\left(-6x^{3}y^{2}+27y^{4}\right)$
$+s^{11}\left(-24xy^{3}+24y^{5}\right)+s^{10}\left(3x^{5}-18x^{2}y^{2}-42xy^{4}\right)$
$+s^{9}\left(12x^{2}y^{3}-12y^{3}\right)+s^{8}\left(15x^{4}+6x^{3}y^{2}-18xy^{2}-42y^{4}\right)$
$+s^{7}\left(24xy^{3}-15y^{5}\right)+s^{6}\left(30x^{3}+18x^{2}y^{2}+15xy^{4}-6y^{2}\right)$
$+12s^{5}y^{3}+s^{4}\left(27x^{2}+18xy^{2}+15y^{4}\right)+3s^{3}y^{5}+s^{2}\left(9x+6y^{2}\right).$
It follows that $s^{2}$ should be canceled out from $\overline{E}(x,y,s)_{1}$. Similarly, one deduces that $s^{3}$ should be canceled out from $\overline{E}(x,y,s)_{2}$. Let $\tilde{E}(x,y,s)$ denote the resulting set of equations. Plug-in $x=[s^{n_{1}+1}]\xi_{1}(s)=\frac{-2i}{3},~y=[s^{n_{2}+1}]\xi_{2}(s)=\frac{\sqrt{2}}{2}+\frac{\sqrt{2}i}{2}$ and $s=0$ into $\tilde{E}$ and observe that the initial conditions hold and that the Jacobian with respect to $x,y$ (evaluated at the initial conditions)
$J=\left[\begin{matrix}9&6\sqrt{2}+6\sqrt{2}i\\ 0&15\left(\frac{\sqrt{2}}{2}+\frac{\sqrt{2}i}{2}\right)^{4}+3\end{matrix}\right]$
is indeed non-singular, concluding the derivation. Although only the first few terms are used, by uniqueness, the rest of the terms obtained by the Newton polyhedron argument are correct.